OpenStudy (dtan5457):

If 4 fair dice are rolled, what is the probability that at least 2 are fives?

OpenStudy (kropot72):

\[\large P(2\ fives)=4C2 \times (\frac{1}{6})^{2} \times(\frac{5}{6})^{2}\] Then we need to use the binomial distribution again to find P(3 fives) and P(4 fives) and add the three values of probability together.

OpenStudy (kropot72):

@dtan5457 Do you follow?